OpenStudy (anonymous):

is (-3,4) a solution of 3x-7y=16?

jimthompson5910 (jim_thompson5910):

3x-7y=16 3(-3)-7(4)=16 -9-28=16 -37=16 ... which is false So (-3,4) is NOT a solution to 3x-7y=16

OpenStudy (anonymous):

aaahhhh I see thanks :D

OpenStudy (anonymous):

substitute the given value of x and y to the equation..and if its not equal with the constant term..it is not the value of x and y..

jimthompson5910 (jim_thompson5910):

yep, just plug in the corresponding values, evaluate/simplify and see if the final equation is true or not